jigglypuff314 (jigglypuff314):

large int_{}^{} \frac{ 2x }{ x^2+6x+13 } Could someone help please? :3 I've learned logs and inverse trig so far :)

OpenStudy (anonymous):

Latex is broken, but seems you just need to get the derivative of the bottom?

jigglypuff314 (jigglypuff314):

\[\large \int\limits_{}^{} \frac{ 2x }{ x^2+6x+13 } \]

jigglypuff314 (jigglypuff314):

derivative of bottom is 2x + 6 ?

OpenStudy (anonymous):

Yeah I'm thinking you have to do a u sub with x^2+6x+13 :d

OpenStudy (anonymous):

Try it out, and see what you get

jigglypuff314 (jigglypuff314):

(du+6)/u ? lol 0.o

ganeshie8 (ganeshie8):

maybe split it into two integrals first

ganeshie8 (ganeshie8):

\(\large \int \frac{2x}{x^2+6x+13} = \int \frac{2x+6-6}{x^2+6x+13} \) \(\large = \int \frac{2x+6}{x^2+6x+13} + \int \frac{-6}{x^2+6x+13}\)

ganeshie8 (ganeshie8):

you could try u-sub for first integral, and a trig-sub for last integral

OpenStudy (mathmale):

The derivative of the denominator is 2x+6, which you unfortately do not have in the numerator. Follow the advice from ganeshie8, or, alternatively, "complete the square" of the denominator.

OpenStudy (anonymous):

Yeah, I just did complete the square, and got the right answer :d

OpenStudy (mathmale):

Cool. guide Jigglypuff towards doing this herself, if she asks you to.

jigglypuff314 (jigglypuff314):

I understand that 1/(x^2+6x+13) -> (1/2)arctan((x+3)/2) but idk what the -6 in the numerator does to that...

OpenStudy (mathmale):

First of all, ganeshie8's approach WILL work. I'll leave it to him/her to explain where those 6's in the numerator came from and why they're helpful.

ganeshie8 (ganeshie8):

\(\large \int c f(x) dx = c\int f(x) dx\)

ganeshie8 (ganeshie8):

just pull out the -6 out of integral

ganeshie8 (ganeshie8):

\(\large \int \frac{2x}{x^2+6x+13} = \int \frac{2x+6-6}{x^2+6x+13} \) \(\large = \int \frac{2x+6}{x^2+6x+13} + \int \frac{-6}{x^2+6x+13}\) \(\large = \int \frac{2x+6}{x^2+6x+13} -6 \int \frac{1}{x^2+6x+13}\)

jigglypuff314 (jigglypuff314):

Oh right ^_^ ok, I'll try that, thank you :)

ganeshie8 (ganeshie8):

np :) did u complete the square for second integral and did trig-sub ?

jigglypuff314 (jigglypuff314):

yep :) I got it, thank you! :D

ganeshie8 (ganeshie8):

okay nice :)

ganeshie8 (ganeshie8):

numerator : x = 2u-3

OpenStudy (phi):

@ganeshie8 good point!